# Show Me What and Where has Changed? Question Answering and Grounding for Remote Sensing Change Detection

### v1

@@ Abstract
Remote sensing change detection aims to perceive changes occurring on the Earth’s surface from remote sensing data in different periods, and feed these changes back to humans. However, most existing methods only focus on detecting change regions, lacking the ability to interact with users to identify changes that the users expect. In this paper, we introduce a new task named Change Detection Question Answering and Grounding (CDQAG), which extends the traditional change detection task by providing interpretable textual answers and intuitive visual evidence. To this end, we construct the first CDQAG benchmark dataset, termed QAG-360K, comprising over 360K triplets of questions, textual answers, and corresponding high-quality visual masks. It encompasses 10 essential land-cover categories and 8 comprehensive question types, which provides a large-scale and diverse dataset for remote sensing applications. Based on this, we present VisTA, a simple yet effective baseline method that unifies the tasks of question answering and grounding by delivering both visual and textual answers. Our method achieves state-of-the-art results on both the classic CDVQA and the proposed CDQAG datasets. Extensive qualitative and quantitative experimental results provide useful insights for the development of better CDQAG models, and we hope that our work can inspire further research in this important yet underexplored direction. The proposed benchmark dataset and method are available at https://github.com/like413/VisTA.
@@ 1 Introduction
Multi-temporal remote sensing imagery plays an essential role in the rapid acquisition of Earth’s surface change information, with applications across various fields such as environmental protection, resource management, and damage assessment [44, 27, 32]. However, existing remote sensing change detection systems primarily focus on full-picture analysis of coverage variations, such as binary segmentation and caption generation. Such systems lack the flexibility to accommodate human instructions [42, 1], thus constraining the advancement of user-friendly and effective remote sensing intelligent interpretation tasks.
Prior to this paper, several studies have attempted to enhance human-computer interaction in change detection tasks through natural language question answering (e.g., CDVQA [41]). However, these methods neglect the intuitive visual explanations, i.e., the change masks referred to by the complex questions, which fail to establish a link between textual answers and the corresponding visual change regions. To this end, we first propose a novel task termed Change Detection Question Answering and Grounding (CDQAG). As illustrated in Fig. 1, this task supports the simultaneous generation of textual answers and their associated pixel-level visual masks, providing an intuitive means for users to verify the answers and increasing their confidence in the reliability of the results. Therefore, exploring how to transcend the limitations of groundless answers and thoroughly investigate the semantic associations between the textual answers and visual groundings is of critical importance and constitutes the core research focus of this paper.
To propel the research of CDQAG, we construct the first benchmark dataset named QAG-360K, by extensively surveying, collecting, and standardizing 3 existing change detection datasets, generating 6,810 image pairs and over 360K high-quality {question, answer, mask} triples through our automated data generation engine. Notably, the change queries include not only straightforward references (e.g., “playground”), but also more complicated description involving complex reasoning (e.g., “the land cover category with the smallest change” or “the area of land cover that increased”). Fig. 1 shows some examples of CDQAG task. For example, given “How much building area has changed in the post-change image?”, the output should provide both the proportion range of “building” and the corresponding change mask. Another example is “What is the smallest change in the second image?”, where the answer necessitates perceiving and evaluating changes across different categories to produce a correct textual-visual answer. Sec. 3 provides a comprehensive description and statistical analysis of the proposed dataset.
@@
To accomplish this task, we develop a powerful benchmark framework called VisTA, which aims to tackle two key challenges: (1) How to reason about complex questions jointly with change images? (2) How to uncover the intrinsic relationships between the cross-modal answers? To solve the first challenge, the Multi-Stage Reasoning Module is designed to adaptively enhance cross-modal information interaction and fusion by jointly leveraging pixel-level visual representations and fine-granularity textual features. Notably, based on the observation that answer grounding typically depend on specific semantic cues in the question or answer, we introduce a question-answer selection module to further refine the reasoning process. While the Text-Visual Answer Decoder is proposed to address the second challenge, which employs a text-to-pixel contrastive learning strategy to achieve fine-grained alignment between textual answers and visual masks.
@@
  • We propose a novel CDQAG task for remote sensing change detection. Unlike classic VQA (see Fig. 1), CDQAG not only generates textual answers but also provides pixel-level visual evidences, which is crucial for developing reliable remote sensing change detection systems.
  • We construct the first CDQAG benchmark dataset QAG-360K, containing over 360K {question, answer, mask} triples that are not only diverse but also contain comprehensive questions. This benchmark is essential for evaluation and encourages the community to further explore the CDQAG task.
  • We present a solid CDQAG baseline method VisTA, which surpasses the latest visual question answering and visual grounding methods, achieving state-of-the-art performance on both the classic CDVQA and QAG-360K datasets.
Figure 2: The proposed QAG-360K benchmark statistics: (a) the answer categories with frequency; (b) the distribution of question types (CN: change or not; CtW: change to what; CfW: change from what; IN: increase or not; DN: decrease or not; LC: largest change; SC: smallest change; CR: change ratio); (c) the word cloud of all questions; (d) the distribution of area ratio for visual masks.
@@ 2 Related work
Visual Question Answering. As a pivotal task in multimodal research, visual question answering (VQA) challenges models to respond to natural language queries based on images, making it a focal point in the domain of vision-language research. In recent years, the release of numerous benchmark datasets, including VQA [8], COCO-QA [24], Visual Genome [15], and CLEVR [13], has significantly accelerated progress in this field. Simultaneously, VQA also demonstrates its enormous potential in many applications such as education, healthcare, and remote sensing [22, 33, 11]. In particular, the introduction of CDVQA [41] has brought VQA into the domain of change detection, establishing a baseline dataset and method that expand the research frontiers of VQA while providing fresh perspectives for multimodal analysis in remote sensing.
Visual Grounding. Visual grounding (VG) focuses on localizing specific objects within images based on natural language descriptions, and is typically divided into two subtasks: referring expression comprehension (REC) [40, 3, 38] and referring expression segmentation (RES) [36, 39]. REC aims to predict the bounding boxes of target objects from linguistic descriptions, while RES requires the generation of pixel-level segmentation masks of the referred targets. Recently, VG in remote sensing [17, 19] has gained increasing attention for its ability to intuitively highlight targets of interest against irrelevant backgrounds.
@@ 3.2 QAG-360K: A Large-scale CDQAG Dataset
To address the gap in this critical yet underexplored area, we introduce QAG-360K, the first benchmark dataset tailored for the CDQAG task. We collected a high-quality set of remote sensing images from existing binary and semantic change detection datasets, including Hi-UCD [29], SECOND [37], and LEVIR-CD [2]. These data cover 24 different regions across cities in Estonia, China and the United States, with spatial resolution ranging from 0.1 to 3.0 meters, providing diverse geographical scenarios. We filtered 6,810 remote sensing image pairs from these datasets as part of QAG-360K, including 10 land-cover categories, all of which are equipped with semantic masks. Additionally, we developed a specialized Triplet Generation Engine (in Appendix) to automatically produce change-related questions, textual answers, and corresponding visual masks.
Question Raised. To encompass a broad range of change scenarios, we expanded the change-related questions into 8C Questions, covering the critical change detection types: change or not? change to what? change from what? increase or not? decrease or not? largest change? smallest change? change ratio? Moreover, to ensure that the phrasing of the questions conforms to natural communication, we utilized the Large Language Model (LLM) to generate an average of 20 question templates for each change type and then manually selected the most reasonable 5 for querying. The final questions average 9.5 words in length, with lengths ranging from 4 to 15 words. Each pair of remote sensing images contains an average of 53 triples. Fig. 2(b) and Fig. 2(c) illustrate the distribution of all question types and the word clouds of land-cover categories, respectively. Detailed definitions and templates for each question type are provided in the Appendix.
Answers Generation. For each question type in the 8C Questions, the answer generation follows different judgment rules. Specifically, existing change detection datasets offer mask annotations, which allow us to automatically generate textual answers and the corresponding pixel-level groundings based on predefined rules. The answer generation rules and specific procedures for each question type are detailed in the Appendix. Fig. 2(d) illustrates the distribution of mask area ratios for each visual answer. CDQAG introduces four critical and distinct attributes that are essential for practical applications:
@@
    Fragmented Spatial Masks. Unlike natural images, which generally exhibit object-level semantic features, certain land-cover types in remote sensing imagery are spread across multiple regions, leading to fragmented and discontinuous segmentation masks. Fig. 3(c3) illustrates an example of proportional change in low vegetation. Moreover, when changes are extremely subtle, it poses a greater challenge for the model to accurately identify and segment them.
@@
These characteristics are pervasive throughout our dataset, ensuring that it maintains a balanced level of complexity. Furthermore, more examples from the QAG-360K dataset are provided in the Appendix.
@@ 4 Method
Next, we provide a detailed description of our proposed CDQAG model, VisTA, which delivers reliable textual answers and precise visual feedback. As shown in Fig. 4, our framework consists of three primary components, which are explained below.
@@ 4.1 Text and Image Feature Extraction
Text Encoder. Given a question $Q\in\mathbb{R}^{L}$, we leverage a CLIP pre-trained Transformer [23] to extract text features $F_{t}\in\mathbb{R}^{L\times C}$, where $L$ denotes the question length and $C$ represents the number of feature channels. The input text sequence is surrounded by two special tokens [SOS] and [EOC], where [EOC] is subsequently activated as a sentence-level representation $F_{s}\in\mathbb{R}^{C}$.
@@ 4.4 Training Objective
Following prior works [20, 32], we employ the standard cross-entropy (CE) loss for textual answer classification and per-pixel binary cross-entropy (BCE) loss for visual answer prediction, denoted as $\mathcal{L}_{txt}$ and $\mathcal{L}_{mask}$, respectively. It is noteworthy that we integrate a text-to-pixel contrastive loss into our training objective to enforce alignment between multimodal answers. Specifically, the text-to-pixel contrastive loss measures the similarity between the textual answer $F_{ta}$ and the pixel-level visual answer $F_{va}$ using a dot product. The contrastive loss $\mathcal{L}_{con}$ is defined as:
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{txt}+\lambda_{2}\mathcal{L}_{mask}+\lambda_{3}\mathcal{L}_{con},$ (11)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the three losses during training.
@@ 5.1 Implementation Details
Evaluation Metrics: Besides the widely-used VQA metrics average accuracy (AA) and overall accuracy (OA), we further introduce mean IoU (mIoU) and overall IoU (oIoU) for CDQAG. Detailed definitions of metrics can be found in the Appendix.
Configurations: We employ the CLIP pre-trained Transformer [23] and ResNet-101 [10] as the text and image encoders for all studies. The input image size and maximum sentence length are set to 512$\times$512 pixels and 15 tokens, respectively. Each multi-head attention layer and feed-forward network uses default settings [6]. For the loss function in Eq. 11, we set $\lambda_{1}$ = 0.2 and $\lambda_{2}$ = $\lambda_{3}$ = 1. In the main results (Tab. 2), the second column represents each type in 8C Questions: (CN: change or not; CtW: change to what; CfW: change from what; IN: increaxse or not; DN: decrease or not; LC: largest change; SC: smallest change; CR: change ratio). Following [34], we train the model for 50 epochs using the AdamW [21] optimizer with an initial learning rate of 1e-4. The learning rate is decreased by a factor of 0.1 at the 35th epoch. We train the model with a batch size of 64 on 4 NVIDIA RTX4090 with 24 GPU VRAM. During inference, we binarize the predicted results using a threshold of 0.35 to obtain the final outputs.
Baseline Methods: Considering that existing methods rarely support both textual and visual answers, we add the same text or mask prediction heads as ours to RES models (e.g., CRIS [34], LAVT [39], CGFormer [28], and ETRIS [36]) and VQA models (e.g., RSVQA [20], CDVQA [41], and SOBA [33]). Moreover, since the aforementioned methods are not specifically designed for change detection, to ensure a fair comparison, we concatenate the two images and feed them into the visual backbone in the same way for all methods (except that marked with symbol $\dagger$).
@@ 5.2 Comparisons with State-of-the-art Methods
Results on CDVQA. We also compare our method with SoTA methods on the two CDVQA test subsets, as illustrated in Tab. 3. Our method consistently outperforms the baselines in both AA and OA metrics. Specifically, on Test 1, our method achieves an OA of 73.1%, signifying a substantial improvement of 10.3% over RSVQA, 7.2% over CDVQA, and 3.9% over SOBA. On Test 2, our approach maintains its advantage with an AA of 66.0% and an OA of 69.5%, further demonstrating significant improvements over existing SoTA approaches. These results underscore the robustness and generalization capabilities of our model, particularly in challenging VQA scenarios. More experimental results are presented in Appendix.
@@
Qualitative results. As depicted in Fig. 5, our model accurately answers the change-related questions and provides intuitive and high-quality visual evidence. More visualizations are given in the Appendix. We also show some failure cases in Fig. 6. In the first row, although the prediction is correct and the visual evidence aligns well with the textual answer, the regions of buildings are discrete units in the ground-truth (GT), while our model generates continuous masks with overly smooth boundaries. This discrepancy underscores the difficulty of accurately capturing the discrete nature of certain objects. In the second case, the color of the playground fades over time, leading to the model mistakenly classify it as a “non-vegetated ground surface” that closely resembles the visual change, resulting in an incorrect prediction. This further requires models to perceive subtle visual cues and comprehend context-sensitive differences for more accurate visual answers.

### v2

@@ Abstract
Remote sensing change detection aims to perceive changes occurring on the Earth’s surface from remote sensing data in different periods, and feed these changes back to humans. However, most existing methods only focus on detecting change regions, lacking the capability to interact with users to identify changes that the users expect. In this paper, we introduce a new task named Change Detection Question Answering and Grounding (CDQAG), which extends the traditional change detection task by providing interpretable textual answers and intuitive visual evidence. To this end, we construct the first CDQAG benchmark dataset, termed QAG-360K, comprising over 360K triplets of questions, textual answers, and corresponding high-quality visual masks. It encompasses 10 essential land-cover categories and 8 comprehensive question types, which provides a valuable and diverse dataset for remote sensing applications. Furthermore, we present VisTA, a simple yet effective baseline method that unifies the tasks of question answering and grounding by delivering both visual and textual answers. Our method achieves state-of-the-art results on both the classic change detection-based visual question answering (CDVQA) and the proposed CDQAG datasets. Extensive qualitative and quantitative experimental results provide useful insights for developing better CDQAG models, and we hope that our work can inspire further research in this important yet underexplored research field. The proposed benchmark dataset and method are available at https://github.com/like413/VisTA.
@@ 1 Introduction
Multi-temporal remote sensing imagery plays an essential role in the rapid acquisition of Earth’s surface change information, with applications across various fields such as environmental protection, resource management, and damage assessment [44, 27, 32]. However, existing remote sensing change detection systems primarily focus on full-picture analysis of coverage variations, such as binary segmentation and caption generation. Such systems lack the flexibility to accommodate human instructions [42, 1], thereby constraining the advancement of user-friendly and effective remote sensing intelligent interpretation tasks.
Prior to this paper, several studies have attempted to enhance human-computer interaction in change detection tasks through natural language question answering (e.g., CDVQA [41]). However, these methods neglect the intuitive visual explanations, i.e., the change masks referred to by the complex questions, which fail to establish a link between textual answers and the corresponding visual change regions. To this end, we propose a novel task termed Change Detection Question Answering and Grounding (CDQAG). As illustrated in Fig. 1, this task supports the simultaneous generation of textual answers and their associated pixel-level visual masks, providing an intuitive means for users to verify the answers and increasing their confidence in the reliability of the results. Therefore, exploring how to transcend the limitations of proofless answers and thoroughly investigate the semantic associations between the textual answers and visual groundings is of critical importance and constitutes the core research focus of this paper.
To propel the research of CDQAG, we construct the first benchmark dataset named QAG-360K, by extensively surveying, collecting, and standardizing 3 existing change detection datasets, generating 6,810 image pairs and over 360K high-quality {question, answer, mask} triples through our automated data generation engine. Notably, the change queries include not only straightforward references (e.g., “playground”), but also more complicated description involving complex reasoning (e.g., “the land cover category with the smallest change” or “the area of land cover that increased”). Fig. 1 shows some examples of CDQAG task. For example, given “How much building area has changed in the post-change image?”, the output should provide both the proportion range of “building” and the corresponding change mask. Another example is “What is the smallest change in the second image?”, where the answer necessitates perceiving and evaluating changes across different categories to generate a correct textual-visual answer. Sec. 3 provides a comprehensive description and statistical analysis of the proposed dataset.
@@
To accomplish this task, we develop a powerful benchmark framework called VisTA, which aims to tackle two key challenges: (1) How to reason about complex questions jointly with change images? (2) How to uncover the intrinsic relationships between the textual answers and the corresponding visual answers? To solve the first challenge, the Multi-Stage Reasoning Module is designed to adaptively enhance cross-modal information interaction and fusion by jointly leveraging pixel-level visual representations and fine-grained textual features. Notably, based on the observation that answer grounding typically depend on specific semantic cues in the question or answer, we introduce a question-answer selection module to further refine the reasoning process. While the Text-Visual Answer Decoder is proposed to address the second challenge, which employs a text-to-pixel contrastive learning strategy to achieve fine-grained alignment between textual answers and visual masks.
@@
  • We propose a novel CDQAG task for remote sensing change detection. Unlike classic VQA (see Fig. 1), CDQAG not only generates textual answers but also provides pixel-level visual evidence, which is crucial for developing reliable remote sensing change detection systems.
  • We construct the first CDQAG benchmark dataset QAG-360K, containing over 360K {question, answer, mask} triples that are not only diverse in land-cover categories but contain comprehensive questions. This benchmark is essential for evaluation and encourages the community to further explore the CDQAG task.
  • We present a solid CDQAG baseline method VisTA, which surpasses the latest visual question answering and visual grounding methods, achieving state-of-the-art performance on both the classic CDVQA and the porposed QAG-360K datasets.
Figure 2: The proposed QAG-360K benchmark statistics: (a) the frequency of answer categories; (b) the distribution of question types (CN: change or not; CtW: change to what; CfW: change from what; IN: increase or not; DN: decrease or not; LC: largest change; SC: smallest change; CR: change ratio); (c) the word cloud of all questions; (d) the distribution of area ratio for visual masks.
@@ 2 Related work
Visual Question Answering. As a pivotal task in multimodal research, visual question answering (VQA) requires models to respond to natural language queries based on images, making it a focal point in the domain of vision-language research. In recent years, the release of numerous benchmark datasets, including VQA [8], COCO-QA [24], Visual Genome [15], and CLEVR [13], has significantly accelerated progress in this field. Simultaneously, VQA also demonstrates its enormous potential in many applications such as education, healthcare, and remote sensing [22, 33, 11]. In particular, the introduction of CDVQA [41] has brought VQA into the domain of change detection, establishing a baseline dataset and method that expand the research frontiers of VQA while providing fresh perspectives for multimodal analysis in remote sensing.
Visual Grounding. Visual grounding (VG) focuses on localizing specific objects within images based on natural language descriptions, and can be typically divided into two subtasks: referring expression comprehension (REC) [40, 3, 38] and referring expression segmentation (RES) [36, 39]. REC aims to predict the bounding boxes of target objects from linguistic descriptions, while RES requires the generation of pixel-level segmentation masks of the referred targets. Recently, VG in remote sensing [17, 19] has gained increasing attention for its ability to intuitively highlight targets of interest against irrelevant backgrounds.
@@ 3.2 QAG-360K: A Large-scale CDQAG Dataset
To address the gap in this critical yet underexplored area, we introduce QAG-360K, the first benchmark dataset tailored for the CDQAG task. We collected a high-quality set of remote sensing images from existing binary and semantic change detection datasets, including Hi-UCD [29], SECOND [37], and LEVIR-CD [2]. These data cover 24 different regions across cities in Estonia, China and the United States, with spatial resolution ranging from 0.1 to 3.0 meters, providing diverse geographical scenarios. We filtered 6,810 remote sensing image pairs from these datasets as part of QAG-360K, including 10 land-cover categories, all of which are equipped with semantic masks. Additionally, we developed a specialized Triplet Generation Engine (in Appendix A) to automatically produce change-related questions, textual answers, and corresponding visual masks.
Question Raised. To encompass a broad range of change scenarios, we expanded the change-related questions into 8C Questions, covering the critical change detection types: change or not? change to what? change from what? increase or not? decrease or not? largest change? smallest change? change ratio? Moreover, to ensure that the phrasing of the questions conforms to natural communication, we utilized the large language model (LLM) to generate an average of 20 question templates for each change type and then manually selected the most reasonable 5 for querying. The final questions average 9.5 words in length, with lengths ranging from 4 to 15 words. Each pair of remote sensing images contains an average of 53 triples. Fig. 2(b) and Fig. 2(c) illustrate the distribution of all question types and the word clouds of land-cover categories, respectively. Detailed definitions and templates for each question type are provided in the Appendix A.2.
Answers Generation. For each question type in the 8C Questions, the answer generation follows different judgment rules. Specifically, existing change detection datasets offer mask annotations, which allow us to automatically generate textual answers and the corresponding pixel-level groundings based on predefined rules. The answer generation rules and specific procedures for each question type are detailed in the Appendix A.2. Fig. 2(d) illustrates the distribution of mask area ratios for each visual answer. CDQAG introduces four critical and distinct attributes that are essential for practical applications:
@@
    Fragmented Spatial Masks. Unlike natural images, which generally exhibit object-level semantic features, certain land-cover categories in remote sensing imagery are spread across multiple regions, leading to fragmented and discontinuous segmentation masks. Fig. 3(c3) illustrates an example of proportional change in low vegetation. Moreover, when changes are extremely subtle, it poses a greater challenge for the model to accurately identify and segment them.
@@
These characteristics are pervasive throughout our dataset, ensuring that it maintains a balanced level of complexity. Furthermore, more examples from the QAG-360K dataset are provided in the Appendix B.2.
@@ 4 Method
Next, we provide a detailed description of our proposed CDQAG model VisTA, which delivers reliable textual answers and precise visual feedback. As shown in Fig. 4, our framework consists of three primary components, which are explained below.
@@ 4.1 Text and Image Feature Extraction
Text Encoder. Given a question $Q\in\mathbb{R}^{L}$, we leverage a CLIP pre-trained Transformer [23] to extract text features $F_{t}\in\mathbb{R}^{L\times C}$, where $L$ denotes the question length and $C$ represents the number of feature channels. The input text sequence is surrounded by two special tokens [SOS] and [EOS], where [EOS] is subsequently activated as a sentence-level representation $F_{s}\in\mathbb{R}^{C}$.
@@ 4.4 Training Objective
Following prior works [20], we employ the standard cross-entropy (CE) loss for textual answer classification, denoted as $\mathcal{L}_{txt}$. For visual evidence, it is noteworthy that we incorporate a text-to-pixel contrastive loss into our training objective to ensure alignment with the corresponding textual answer. Specifically, the text-to-pixel contrastive loss measures the similarity between the textual answer $F_{ta}$ and the pixel-level visual answer $F_{va}$ using a dot product. The contrastive loss $\mathcal{L}_{con}$ is defined as:
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{txt}+\lambda_{2}\mathcal{L}_{con},$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that balance the three losses during training.
@@ 5.1 Implementation Details
Evaluation Metrics: Besides the widely-used VQA metrics average accuracy (AA) and overall accuracy (OA), we further introduce mean IoU (mIoU) and overall IoU (oIoU) for CDQAG. Detailed definitions of metrics can be found in the Appendix C.1.
Configurations: We employ the CLIP pre-trained Transformer [23] and ResNet-101 [10] as the text and image encoders for all studies. The input image size and maximum sentence length are set to 512$\times$512 pixels and 40 tokens, respectively. Each multi-head attention layer and feed-forward network uses default settings [6]. For the loss function in Eq. 11, we set $\lambda_{1}$ = 0.2 and $\lambda_{2}$ = 1. In the main results (Tab. 2), the second column represents each type in 8C Questions: (CN: change or not; CtW: change to what; CfW: change from what; IN: increaxse or not; DN: decrease or not; LC: largest change; SC: smallest change; CR: change ratio). Following [34], we train the model for 50 epochs using the AdamW [21] optimizer with an initial learning rate of 1e-4. The learning rate is decreased by a factor of 0.1 at the 35th epoch. We train the model with a batch size of 64 on 4 NVIDIA RTX4090 with 24 GPU VRAM. During inference, we binarize the predicted results using a threshold of 0.35 to obtain the final outputs.
Baseline Methods: Considering that existing methods rarely provide both textual and visual answers, we add the same text or mask prediction heads as ours to RES models (e.g., CRIS [34], LAVT [39], CGFormer [28], and ETRIS [36]) and VQA models (e.g., RSVQA [20], CDVQA [41], and SOBA [33]). Moreover, since the aforementioned methods are not specifically designed for change detection, to ensure a fair comparison, we concatenate the two images and feed them into the visual backbone in the same way for all methods (except that marked with symbol $\dagger$).
@@ 5.2 Comparisons with State-of-the-art Methods
Results on CDVQA. We also compare our method with SoTA methods on the two CDVQA test subsets, as illustrated in Tab. 3. Our method consistently outperforms the baselines in both AA and OA metrics. Specifically, on Test 1, our method achieves an OA of 73.1%, signifying a substantial improvement of 10.3% over RSVQA, 7.2% over CDVQA, and 3.9% over SOBA. On Test 2, our approach maintains its advantage with an AA of 66.0% and an OA of 69.5%, further demonstrating significant improvements over existing SoTA approaches. These results underscore the robustness and generalization capabilities of our model, particularly in challenging VQA scenarios. More experimental results are presented in Appendix C.3.
@@
Qualitative results. As depicted in Fig. 5, our model accurately answers the change-related questions and provides intuitive and high-quality visual evidence. More visualizations are given in the Appendix C.4. We also show some failure cases in Fig. 6. In the first row, although the prediction is correct and the visual evidence aligns well with the textual answer, the regions of buildings are discrete units in the ground-truth (GT), while our model generates continuous masks with overly smooth boundaries. This discrepancy underscores the difficulty of accurately capturing the discrete nature of certain objects. In the second case, the color of the playground fades over time, leading to the model mistakenly classify it as a “non-vegetated ground surface” that closely resembles the visual change, resulting in an incorrect prediction. This further requires models to perceive subtle visual cues and comprehend context-sensitive differences for more accurate visual answers.